# MiniGPT-5: Interleaved Vision-and-Language Generation via Generative Vokens

### v3

@@ Abstract
The effectiveness of Multimodal Large Language Models (MLLMs) demonstrates a profound capability in multimodal understanding. However, the simultaneous generation of images with coherent texts is still underdeveloped. Addressing this, we introduce a novel interleaved vision-and-language generation method, centered around the concept of “generative vokens”. These vokens serve as pivotal elements contributing to coherent image-text outputs. Our method is marked by a unique two-stage training strategy for description-free multimodal generation, which does not necessitate extensive descriptions of images. We integrate classifier-free guidance to enhance the alignment of generated images and texts, ensuring more seamless and contextually relevant multimodal interactions. Our model, MiniGPT-5, exhibits substantial improvement over the baseline models on multimodal generation datasets, including MMDialog and VIST. The human evaluation shows MiniGPT-5 is better than the baseline model on more than 56% cases for multimodal generation, highlighting its efficacy across diverse benchmarks.
@@ 1 Introduction
The development of large-scale vision-and-language models is significantly impacting a wide range of fields like automated dialogue systems and digital content creation. With the surge in research and development in this domain, the current state-of-the-art Large Language Models (LLMs) (OpenAI, 2023; Chiang et al., 2023; Ouyang et al., 2022) and vision-and-language models such as (Wu et al., 2023a; Li et al., 2023c; Tsimpoukelli et al., 2021; Alayrac et al., 2022) fall short in generating coherent multimodal outputs. This limitation becomes particularly evident in tasks that demand an integrated handling of vision and language, essential for the next generation Large Language Models (LLMs).
@@
Our work, as illustrated in Fig. 1, seeks to address these shortcomings by enhancing the integration of text and image generation in LLMs. The challenges in developing a multimodal LLM capable of interleaved vision and language generation are manifold. First, LLMs typically lack mechanisms to directly produce images, prompting us to introduce “generative vokens” that bridge the gap between textual and visual feature spaces. Second, the constraint of data scarcity, especially in vision-and-language tasks (Sharma et al., 2018) lacking extensive detailed descriptions of images (Huang et al., 2016), is countered by our unique description-free training approach. Third, maintaining both image-text and image-image consistency poses a significant challenge, which we address through dual-loss strategies. Finally, as we push forward the boundaries with LLMs, the large memory requirements urge us to devise more efficient end-to-end strategies and create an efficient training pipeline accessible for the community, especially in downstream tasks.
@@
Specifically, to overcome these challenges, we present MiniGPT-5, a novel approach for interleaved vision-and-language generation. By combing the Stable Diffusion with LLMs through special visual tokens (Tan & Bansal, 2020) – “generative vokens”, we develop a new approach for multimodal generation. Our two-stage training methodology emphasizes a description-free foundational phase, enabling effective model training even with limited caption-grounded images. This strategy, distinct from existing works, pivots on generic stages free from image annotations. To ensure that the generated text and images are in harmony, our dual-loss strategy comes into play, further enhanced by our innovative generative voken approach and classifier-free guidance. Our parameter-efficient fine-tuning strategy optimizes training efficiency and addresses memory constraints.
@@
As shown in Fig. 2, leveraging ViT (Vision Transformer) and Qformer (Li et al., 2023c), alongside Large Language Models, we adapt multimodal inputs into generative vokens, seamlessly combined with the high-resolution Stable Diffusion 2.1 model (Rombach et al., 2022b) for context-aware image generation. Incorporating images as auxiliary input with instruction tuning approaches and pioneering both the text and image generation loss, we amplify the synergy between text and visuals. We experiment on the CC3M (Sharma et al., 2018), VIST (Huang et al., 2016), and MMDialog (Feng et al., 2022) datasets. Notably, MiniGPT-5 shows superior performance across the two multimodal generation datasets.
@@ 2 Related Work
Large Language Models As Large Language Models (LLMs) become increasingly impactful and accessible, a growing body of research has emerged to extend these pretrained LLMs into the realm of multimodal comprehension tasks (Zhu et al., 2023; Li et al., 2023c; Dai et al., 2023; OpenAI, 2023; Li et al., 2023a; Alayrac et al., 2022; Li et al., 2023b). For example, to reproduce the impressive multimodal comprehension ability in GPT-4 (OpenAI, 2023), MiniGPT-4 (Zhu et al., 2023) proposes a projection layer to align pretrained vision component of BLIP-2 (Li et al., 2023c) with an advanced open-source large language model, Vicuna (Chiang et al., 2023). In our work, we utilize the MiniGPT-4 as the base model and extend the model’s capabilities to multimodal generation.
Text-to-Image Generation  To transform textual descriptions into their corresponding visual representations, text-to-image models (Reed et al., 2016; Dhariwal & Nichol, 2021; Saharia et al., 2022; Rombach et al., 2022b; a; Gu et al., 2023; Nichol et al., 2021; Ramesh et al., 2021; Yu et al., 2022; Chang et al., 2023) employ complex architectures and sophisticated algorithms, bridging the gap between textual information and visual content. These models are adept at interpreting the semantics of input text and translating them into coherent and pertinent images. A notable recent contribution in this field is Stable Diffusion V2 (Rombach et al., 2022b), which employs a diffusion process to generate conditional image features and subsequently reconstructs images from these features. Our research aims to leverage this pretrained model, enhancing its capabilities to accommodate both multimodal input and output.
Multimodal Generation with Large Language Models  To augment the LLM’s capabilities in seamlessly integrating vision and language generation, recent studies have introduced a variety of innovative methods (Ge et al., 2023; Sun et al., 2021; Koh et al., 2023; Sun et al., 2023b; Yu et al., 2023; Aiello et al., 2023; Wu et al., 2023c). For instance, CM3Leon (Yu et al., 2023) presents a retrieval-augmented, decoder-only architecture designed for both text-to-image and image-to-text applications. Similarly, Emu (Sun et al., 2023b) employs the pretrained EVA-CLIP (Sun et al., 2023a) model to convert images into one-dimensional features and fine-tunes the LLAMA (Touvron et al., 2023) model to generate cohesive text and image features through autoregressive techniques. On the other hand, NextGPT (Wu et al., 2023c), GILL (Koh et al., 2023) and SEED (Ge et al., 2023) explore the concept of mapping vokens into the text feature space of a pretrained Stable Diffusion model; GILL and NextGPT employ an encoder-decoder framework, while SEED utilizes a trainable Q-Former structure. In contrast to these approaches, our model takes a more direct route by aligning voken features with visual information. Additionally, we introduce several training strategies to enhance image quality and contextual coherence.
@@ 3 Method
In order to endow Large Language Models with multimodal generation capabilities, we introduce a new framework that integrates pretrained multimodal Large Language Models and text-to-image generation models. Central to our approach is the introduction of “generative vokens”, special visual tokens that effectively bridge the textual and visual domains during the training process. Additionally, we implement a two-stage training method combined with a classifier-free guidance strategy to enhance the quality and coherence of generated outputs. Fig. 2 provides an overview of our model structure. MiniGPT-5 primarily consists of two modules: the Integrated Vision-Language Encoding Module, utilizing the pretrained multimodal large language model (MiniGPT-4) for handling multimodal inputs, and the Multimodal Output Generation module, employing Stable Diffusion for generating visual outputs.
@@ 3.1 Multimodal Understanding Module
Recent advancements in multimodal Large Language Models, such as MiniGPT-4 (Zhu et al., 2023), have primarily concentrated on multimodal comprehension, enabling the processing of images as sequential input. The Integrated Vision-Language Encoding Module is designed to extend the capabilities of LLMs from mere comprehension to active generation in multimodal contexts. Generative vokens play a crucial role in this module, enabling the translation of raw visual inputs into a format that LLMs can process and utilize for subsequent generation tasks.
@@
[…]
@@
Parameter-Efficient Fine-Tuning (PEFT) Parameter-efficient fine-tuning (PEFT) (Houlsby et al., 2019; Hu et al., 2021; Li & Liang, 2021) is critical in training Large Language Models (LLMs), employed to adapt LLMs to downstream tasks without the need for extensive retraining. In PEFT, rather than updating all the parameters of a model, only a small subset of parameters is trained. This subset typically includes task-specific components or lightweight layers added to the original model architecture (Zhang et al., 2021; Houlsby et al., 2019; Hu et al., 2021; Dettmers et al., 2023). We apply PEFT to the MiniGPT-4 (Zhu et al., 2023) encoder, enhancing its ability to process and generate multimodal content based on given instructions or prompts. More specifically, this involves the use of prefix tuning (Li & Liang, 2021) and LoRAHu et al. (2021) over the entire language encoder – Vicuna (Chiang et al., 2023) used in MiniGPT-4. Additionally, we implement learnable queries at the input of the transformer decoder, a conventional approach in sequence-to-sequence transformer architectures, to further improve the model’s multimodal generation capabilities. We also adopted learnable queries at the input of the transformer decoder as a conventional setting for sequence-to-sequence transformer architectures (Vaswani et al., 2017a). Learnable queries in the decoder allow the model to have dynamic, adaptable representations for initiating the generation process. This is particularly useful when the model needs to generate outputs based on a mix of visual and textual inputs. Combined with the instruction tuning (Ouyang et al., 2022), it notably amplifies multimodal generation performance across various datasets.
3.2 Mutimodal Generation Module
@@
To accurately align the generative vokens with the text-to-image generation models, we formulate a compact mapping module for dimension matching and incorporate several supervised losses, including voken positioning loss and voken alignment loss. The voken positioning loss assists the model in learning the correct positioning of tokens, while the voken alignment loss directly aligns the vokens with the appropriate conditional generation features of the diffusion model. Since the gradients of generative vokens’ features can be directly calculated from images, shown on the right side of Fig. 2, our method does not need comprehensive descriptions of images, leading to description-free learning.
@@
Voken Positioning  We first jointly generate both text and vokens in the text space by following next-word prediction in autoregressive language model (Vaswani et al., 2017b). During the training, we append the vokens $V_{\text{{img}}}$ to the positions of ground truth images and train the model to predict vokens within text generation. Specifically, the generated tokens are represented as $W=\{w_{1},w_{2},\ldots,w_{m}\}$, where $w_{i}\in V\cup V_{\text{{img}}}$, and the causal language modeling loss is defined as:
@@
To generate appropriate images, the mapping feature $\hat{h}_{\text{{voken}}}$ is used as a conditional input in the denoising process. Intuitively, $\hat{h}_{\text{{voken}}}$ should represent the corresponding conditional features that conduct the diffusion model to generate the ground truth image. We employ the latent diffusion model (LDM) loss as voken alignment loss for training the image generation module. During the training, the ground truth image is first converted to latent feature $z_{0}$ through the pretrained VAE (Variational Autoencoder) (Kingma & Welling, 2013). Then, we obtain the noisy latent feature $z_{t}$ by adding noise $\epsilon$ to $z_{0}$. A pretrained U-Net model $\epsilon_{\theta}$ is used to calculate the conditional LDM loss as:
@@
To summarize, the voken positioning loss enables the model to learn the accurate placement of tokens. Without this component, the model lacks the essential capability to predict when vokens should be generated during inference. Additionally, the voken alignment loss ensures the direct correspondence between vokens and the appropriate conditional generation characteristics of the diffusion model. In the absence of this loss, the model is unable to learn semantic vokens from images directly. This comprehensive approach ensures a coherent understanding and generation of both textual and visual elements, leveraging the capabilities of pretrained models, specialized tokens, and innovative training techniques.
@@ 3.3 Training Strategy
Given the non-negligible domain shift between text and image domains, we observe that direct training on a limited interleaved text-and-image dataset can result in misaligning generated texts and images and diminished image quality. Consequently, we adopt a two-stage training strategy: an initial pretraining stage focusing on coarse feature alignment for unimodal generation, followed by a fine-tuning stage dedicated to intricate feature learning for multimodal generation. Furthermore, to amplify the effectiveness of the generative tokens throughout the diffusion process, we incorporate the idea of classifier-free guidance (Ho & Salimans, 2022) technique through the whole training process.
@@
After the pretraining stage, the model is capable of generating images for single text descriptions but struggles with interleaved vision-and-language generation, which includes multiple text-image pairs and requires complicated reasoning for both text and image generation. To address this, in the fine-tuning stage, we further fine-tune our model with PEFT parameters by interleaved vision-and-language datasets, such as VIST, where the data sample has several steps with text-image and texts are sequentially relevant. During this stage, we construct three types of tasks from the dataset, encompassing (1) text-only generation: given the next image, generating the related text; (2) image-only generation: given the next text, generating the related image, and (3) multimodal generation: generating text-image pair by given context. The fine-tuning stage loss is given by $L_{\text{{Fine-tune}}}=\lambda_{1}*L_{\text{{text}}}+\lambda_{2}*L_{\text{{LDM}}}$. More implementation details can be found in Appendix A.
@@ 4 Experiments
To assess the efficacy of our model, we conducted a series of evaluations across multiple benchmarks. These experiments aim to address several key questions: (1) Can our model generate plausible images and reasonable texts? (2) How does our model compare with state-of-the-art models in both single-turn and multi-turn interleaved vision-and-language generation tasks? (3) What impact does the design of each module have on overall performance? Below we will discuss the experimental setup and present a comprehensive analysis of our model’s performance. We use three datasets: CC3M (Sharma et al., 2018), VIST (Huang et al., 2016), and MMDialog (Feng et al., 2022). More details about datasets and data format can be found in Appendix B.

### v4

# Interleaved Vision-and-Language Generation via Generative Voken

@@ Abstract
The effectiveness of Multimodal Large Language Models (MLLMs) demonstrates a profound capability in multimodal understanding. However, the simultaneous generation of images with coherent texts is still underdeveloped. Addressing this, we introduce a novel interleaved vision-and-language generation method, centered around the concept of “generative vokens”. These vokens serve as pivotal elements contributing to coherent image-text outputs. Our method is marked by a unique two-stage training strategy for description-free multimodal generation, which does not necessitate extensive descriptions of images. We integrate classifier-free guidance to enhance the alignment of generated images and texts, ensuring more seamless and contextually relevant multimodal interactions. Our model, MiniGPT-5, exhibits substantial improvement over the baseline models on multimodal generation datasets, including MMDialog and VIST. The human evaluation shows MiniGPT-5 is better than the baseline model on more than 56% cases for multimodal generation, highlighting its efficacy across diverse benchmarks. Project page: https://eric-ai-lab.github.io/minigpt-5.github.io/.
@@ 1 Introduction
The development of large-scale vision-and-language models is significantly impacting a wide range of fields like automated dialogue systems and digital content creation. With the surge in research and development in this domain, the current state-of-the-art Large Language Models (LLMs) [openai2023gpt4, vicuna2023, training_lm] and vision-and-language models such as [visualchatgpt, li2023blip, frozen, alayrac2022flamingo] fall short in generating coherent multimodal outputs. This limitation becomes particularly evident in tasks that demand an integrated handling of vision and language, essential for the next generation Large Language Models (LLMs).
@@
Our work, as illustrated in Fig. 1, seeks to address these shortcomings by enhancing the integration of text and image generation in LLMs. The challenges in developing a multimodal LLM capable of interleaved vision and language generation are manifold. First, LLMs typically lack mechanisms to directly produce images, prompting us to introduce “generative vokens” that bridge the gap between textual and visual feature spaces. Second, the constraint of data scarcity, especially in vision-and-language tasks [sharma2018conceptual] lacking extensive detailed captions of images [huang2016visual], like descriptions of the elements inside images, is countered by our unique description-free training approach. Third, maintaining both image-text and image-image consistency poses a significant challenge, which we address through dual-loss strategies. Finally, as we push forward the boundaries with LLMs, the large memory requirements urge us to devise more efficient end-to-end strategies and create an efficient training pipeline accessible for the community, especially in downstream tasks.
@@
Specifically, to overcome these challenges, we present MiniGPT-5, a novel approach for interleaved vision-and-language generation. By combining Stable Diffusion with LLMs through special visual tokens [tan2020vokenization] – “generative vokens”, we develop a new approach for multimodal generation. Our two-stage training methodology emphasizes a description-free foundational phase, enabling effective model training even with limited caption-grounded images. This strategy, distinct from existing works, pivots on generic stages free from image annotations. To ensure that the generated text and images are in harmony, our dual-loss strategy comes into play, further enhanced by our innovative generative voken approach and classifier-free guidance. Our parameter-efficient fine-tuning strategy optimizes training efficiency and addresses memory constraints.
@@
As shown in Fig. 2, leveraging ViT (Vision Transformer) and Qformer [li2023blip], alongside Large Language Models, we adapt multimodal inputs into generative vokens, seamlessly combined with the high-resolution Stable Diffusion 2.1 model [rombach2021highresolution] for context-aware image generation. Incorporating images as auxiliary input with instruction tuning approaches and pioneering both the text and image generation loss, we amplify the synergy between text and visuals. We experiment on the CC3M [sharma2018conceptual], VIST [huang2016visual], and MMDialog [feng2022mmdialog] datasets. Notably, MiniGPT-5 shows superior performance across the two multimodal generation datasets.
@@ 2 Related Work
Multimodal Large Language Models As Large Language Models (LLMs) become increasingly impactful and accessible, a growing body of research has emerged to extend these pretrained LLMs into the realm of multimodal comprehension tasks [zhu2023minigpt, li2023blip, instructblip, openai2023gpt4, li2023otter, alayrac2022flamingo, li-etal-2023-lavis, bai2025qwen2]. For example, to reproduce the impressive multimodal comprehension ability in GPT-4 [openai2023gpt4], MiniGPT-4 [zhu2023minigpt] proposes a projection layer to align pretrained vision component of BLIP-2 [li2023blip] with an advanced open-source large language model, Vicuna [vicuna2023]. In our work, we utilize the MiniGPT-4 as the base model and extend the model’s capabilities to multimodal generation.
Text-to-Image Generation  To transform textual descriptions into their corresponding visual representations, text-to-image models [reed2016generative, dhariwal2021diffusion, saharia2022photorealistic, rombach2021highresolution, rombach2022high, gu2023photoswap, nichol2021glide, ramesh2021zero, yu2022scaling, chang2023muse, esser2024scaling] design algorithms to bridge the gap between textual information and visual content. A notable recent contribution is Stable Diffusion V2 [rombach2021highresolution], which employs a diffusion process to generate conditional image features and subsequently reconstructs images from these features. Our research aims to leverage this pretrained model, enhancing its capabilities to accommodate both multimodal input and output.
[…]
@@ 3 Method
In order to endow Large Language Models with multimodal generation capabilities, we introduce a new framework that integrates pretrained multimodal Large Language Models and text-to-image generation models. Central to our approach is the introduction of “generative vokens”, special visual tokens that effectively bridge the textual and visual domains during the training process. Additionally, we implement a two-stage training method combined with a classifier-free guidance strategy to enhance the quality and coherence of generated outputs. Fig. 2 provides an overview of our model structure. MiniGPT-5 primarily consists of two modules: the Multimodal Understanding module, utilizing the pretrained multimodal large language model (MiniGPT-4) for handling multimodal inputs, and the Multimodal Generation module, employing Stable Diffusion for generating visual outputs.
@@ 3.1 Multimodal Understanding Module
Recent advancements in multimodal Large Language Models, such as MiniGPT-4 [zhu2023minigpt], have primarily concentrated on multimodal comprehension, enabling the processing of images as sequential input. The Integrated Vision-Language Encoding Module is designed to extend the capabilities of LLMs from mere comprehension to active generation in multimodal contexts. Generative vokens play a crucial role in this module, enabling the translation of raw visual inputs into a format that LLMs can process and utilize for subsequent generation tasks.
@@
Generative Vokens  Since the original LLM’s $V$ vocabulary only includes the textual tokens, we need to construct a bridge between the LLM and the generative model. Therefore, we introduce a set of special tokens $V_{\text{{img}}}=\{[\text{{IMG1}}],[\text{{IMG2}}],\ldots,[\text{{IMGn}}]\}$ (by default $n=8$) as generative vokens into the LLM’s vocabulary $V$. The LLM’s output hidden state for these vokens is harnessed for subsequent image generation, and the positions of these vokens can represent the insertion of the interleaved images. With all pretrained weights $\theta_{\text{{pretrained}}}$ in MiniGPT-4 fixed, the trainable parameters include extra input embedding $\theta_{\text{{voken\_input}}}$ and output embedding $\theta_{\text{{voken\_output}}}$.
@@
Parameter-Efficient Fine-Tuning (PEFT) Parameter-efficient fine-tuning (PEFT) [houlsby2019parameter, hu2021lora, li2021prefix] is critical in training Large Language Models (LLMs), employed to adapt LLMs to downstream tasks without the need for extensive retraining. In PEFT, rather than updating all the parameters of a model, only a small subset of parameters is trained. This subset typically includes task-specific components or lightweight layers added to the original model architecture [tip_adapter, houlsby2019parameter, hu2021lora, dettmers2023qlora]. We apply PEFT to the MiniGPT-4 [zhu2023minigpt] encoder, enhancing its ability to process and generate multimodal content based on given instructions or prompts. More specifically, this involves the use of prefix tuning [li2021prefix] and LoRA[hu2021lora] over the entire language encoder – Vicuna [vicuna2023] used in MiniGPT-4. Additionally, we implement learnable queries at the input of the transformer decoder, a conventional approach in sequence-to-sequence transformer architectures, to further improve the model’s multimodal generation capabilities. We also adopted learnable queries at the input of the transformer decoder as a conventional setting for sequence-to-sequence transformer architectures [transformer]. Learnable queries in the decoder allow the model to have dynamic, adaptable representations for initiating the generation process. This is particularly useful when the model needs to generate outputs based on a mix of visual and textual inputs. Combined with the instruction tuning [training_lm], it notably amplifies multimodal generation performance across various datasets.
3.2 Multimodal Generation Module
@@
To accurately align the generative vokens with the text-to-image generation models, we formulate a compact mapping module for dimension matching and incorporate several supervised losses, including causal language modeling loss and voken alignment loss. The causal language modeling loss assists the model in learning the correct positioning of tokens, while the voken alignment loss directly aligns the vokens with the appropriate conditional generation features of the diffusion model. Since the gradients of generative vokens’ features can be directly calculated from images, shown on the right side of Fig. 2, our method does not need comprehensive descriptions of images, leading to description-free learning.
@@
Voken Positioning  We first jointly generate both text and vokens in the text space by following next-word prediction in autoregressive language model [vaswani2017attention]. During the training, we append the vokens $V_{\text{{img}}}$ to the positions of ground truth images and train the model to predict vokens within text generation. Specifically, the generated tokens are represented as $W=\{w_{1},w_{2},\ldots,w_{m}\}$, where $w_{i}\in V\cup V_{\text{{img}}}$, and the causal language modeling loss is defined as:
@@
To generate appropriate images, the mapping feature $\hat{h}_{\text{{voken}}}$ is used as a conditional input in the denoising process. Intuitively, $\hat{h}_{\text{{voken}}}$ should represent the corresponding conditional features that conduct the diffusion model to generate the ground truth image. We employ the latent diffusion model (LDM) loss as voken alignment loss for training the image generation module. During the training, the ground truth image is first converted to latent feature $z_{0}$ through the pretrained VAE (Variational Autoencoder) [vae]. Then, we obtain the noisy latent feature $z_{t}$ by adding noise $\epsilon$ to $z_{0}$. A pretrained U-Net model $\epsilon_{\theta}$ is used to calculate the conditional LDM loss as:
@@
To summarize, the causal language modeling loss enables the model to learn the accurate placement of tokens. Without this component, the model lacks the essential capability to predict when vokens should be generated during inference. Additionally, the voken alignment loss ensures the direct correspondence between vokens and the appropriate conditional generation characteristics of the diffusion model. In the absence of this loss, the model is unable to learn semantic vokens from images directly. This comprehensive approach ensures a coherent understanding and generation of both textual and visual elements, leveraging the capabilities of pretrained models, specialized tokens, and innovative training techniques.
@@ 3.3 Training Strategy
Given the non-negligible domain shift between text and image domains, we observe that direct training on a limited interleaved text-and-image dataset can result in misaligning generated texts and images and diminished image quality. Consequently, we adopt a two-stage training strategy: an initial pretraining stage focusing on coarse feature alignment for unimodal generation, followed by a fine-tuning stage dedicated to intricate feature learning for multimodal generation. Furthermore, to amplify the effectiveness of the generative tokens throughout the diffusion process, we incorporate the idea of classifier-free guidance [classifier_free_guidance] technique through the whole training process.
@@
After the pretraining stage, the model is capable of generating images for single text descriptions but struggles with interleaved vision-and-language generation, which includes multiple text-image pairs and requires complicated reasoning for both text and image generation. To address this, in the fine-tuning stage, we further fine-tune our model with PEFT parameters by interleaved vision-and-language datasets, such as VIST, where the data sample has several steps with text-image and texts are sequentially relevant. During this stage, we construct three types of tasks from the dataset, encompassing (1) text-only generation: given the next image, generating the related text; (2) image-only generation: given the next text, generating the related image, and (3) multimodal generation: generating text-image pair by given context. The fine-tuning stage loss is given by $L_{\text{{Fine-tune}}}=\lambda_{1}*L_{\text{{text}}}+\lambda_{2}*L_{\text{{LDM}}}$. More implementation details can be found in Appendix 4.1.
@@ 4 Experiments
To assess the efficacy of our model, we conducted a series of evaluations across multiple benchmarks. These experiments aim to address several key questions: (1) Can our model generate plausible images and reasonable texts? (2) How does our model compare with state-of-the-art models in both single-turn and multi-turn interleaved vision-and-language generation tasks? (3) What impact does the design of each module have on overall performance? Below we will discuss the experimental setup and present a comprehensive analysis of our model’s performance. We use three datasets: CC3M [sharma2018conceptual], VIST [huang2016visual], and MMDialog [feng2022mmdialog]. More details about datasets and data format can be found in Appendix A.
4.1 Implementation Details